# Spontaneous Symmetry Breaking in Graviweak Theory

### v1

@@ I Introduction and Motivation
The unification of gravity with the other fundamental interactions remains one of the most profound challenges in theoretical physics. While the Standard Model successfully unifies the electroweak and strong interactions within a gauge-theoretic framework, gravity remains conceptually and technically distinct due to its geometric nature. An intriguing approach to unification is the idea that gravity and the weak interaction could emerge from a common underlying chiral gauge theory. This framework is referred to as graviweak unification Alexander:2012ge ; Alexander:0706.4481 ; Nesti:0706.3307 ; Nesti:0706.3304 . The motivation stems from the realization that both gravity and the weak interaction can be formulated as chiral theories. The weak interaction is intrinsically chiral—only left-handed fermions couple to the weak gauge bosons—while gravity, in the Einstein-Cartan formalism, can also be made chiral by splitting the spin connection into left- and right-handed components associated with the group structure KirillBook ; Freidel:2005mq ; Capovilla:1991qt ; Plebanski:1977pw ; Celada:2016jdt ; Alexandre:2407.19363
@@
These considerations motivate the construction of a theory in which gravity is initially described by both chiral components. A spontaneous symmetry-breaking mechanism then suppresses the right-handed chiral component at low energies. The surviving left-handed component would naturally couple to left-handed fermions, mimicking the chiral nature of the weak interaction. In particular, the gravitational interaction should have the opposite chirality of the weak interaction, reflecting the parity-breaking structure of the Standard Model. This paper develops a model of graviweak unification where a Higgs-like field higgs generates this parity asymmetry dynamically through a $Z_{2}$ symmetric potential. The right-handed kinetic term becomes heavy and decouples, leaving a chiral gravitational sector.
We develop these ideas in this paper along the following logical thread. In Section II we propose a gravitational theory formulated in terms of chiral components of the Lorentz connection and tetrad fields. The total Lagrangian includes Einstein-Hilbert-type terms, Yang-Mills-like kinetic terms for the chiral curvatures, and a Higgs-like antisymmetric tensor field $\phi^{ab}$ that couples to the chiral sectors and dynamically breaks parity symmetry through its vacuum expectation value. Thus, we have a theory that unifies gravity and the weak interaction in a parity invariant construction before symmetry breaking, reducing to the Ashtekar self-dual formulation of gravity plus the weak sector after symmetry breaking. This will be the foundations of this paper, around which we will investigate developments and variations.
In Section III we then rewrite this theory in the two-spinor notation, the natural natural language for treating gravity and chiral gauge fields on the same footing. This formalism is especially convenient for formulating unification frameworks such as graviweak models and it points us towards a new development: parity symmetry breaking that goes hand in hand with a larger gauge group. This receives further exploration in Section IV, where we extend the internal Lorentz group underpinning the basic construction to $SO(1,3+N)$. A major implication is the appearance of possible interactions between the gravitational 2-form curvature $R^{a}_{~{}b}$ of $SO(1,3)$ and the Yang-Mills field strength $F^{ij}$ of a $SO(N)$ group. This is a remarkable novelty with regards to previous graviweak models. We close this paper with a few speculations on the testability and falsifiability of this scenario.
@@ II The Graviweak-Higgs System
We start by laying down the structure of the theory. The theory is initially parity invariant and built from left- and right-handed chiral curvature 2-forms, defined as:
$R^{(\pm)}_{ab}=d\omega^{(\pm)}_{ab}+\omega^{(\pm)}_{a}{}^{c}\wedge\omega^{(\pm)}_{cb}$ (1)
@@
$\omega^{(\pm)}_{ab}=\frac{1}{2}\left(\omega_{ab}\pm\frac{i}{2}\epsilon_{ab}{}^{cd}\omega_{cd}\right).$ (2)
@@
We then consider the most general parity-invariant gauge theory containing only linear and quadratic terms in the curvature, with the condition that the theory should result in general relativity and the weak sector after spontaneous symmetry breaking. Its Lagrangian is:
@@
$\mathcal{L}_{\text{total}}=\;$ $\frac{1}{32\pi G}\,\epsilon^{abcd}\,e_{a}\wedge e_{b}\wedge\left(R^{(+)}_{cd}+R^{(-)}_{cd}\right)$
$+\frac{1}{4g_{+}^{2}}R^{(+)}_{ab}\wedge\star R^{ab(+)}+\frac{1}{4g_{-}^{2}}R^{(-)}_{ab}\wedge\star R^{ab(-)}$
$+\frac{\alpha}{2}\,\phi^{bc}\left(R^{(+)}_{ab}\wedge\star R^{a(+)}_{~{}c}+R^{(-)}_{ab}\wedge\star R^{a(-)}_{~{}c}\right)$
$-\frac{1}{2}D\phi^{ab}\wedge*D\phi_{ab}+\frac{V(\phi)}{24}\epsilon^{abcd}e_{a}\wedge e_{b}\wedge e_{c}\wedge e_{d}$ (3)
where the star symbol $\star$ denotes the Hodge dual on the frame indices and the asterisk * denotes the Hodge dual on the spacetime indices:
$*R^{ab}_{\mathrm{\ \ }\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}R^{ab\rho\sigma}$ (4)
and
$\star R^{ab}_{\mathrm{\ \ }\mu\nu}=\frac{1}{2}\epsilon^{abcd}R_{cd\mu\nu}.$ (5)
@@
A few comments on this action are in order. First note that the parity transformations under which this theory is initially symmetric are defined as follows. The tetrad transforms as $e^{a}(x^{0},\vec{x})\mapsto e^{a}(x^{0},-\vec{x})$, and the chiral components of the connection exchange under parity: $\omega^{(+)}_{ab}(x^{0},\vec{x})\mapsto\omega^{(-)}_{ab}(x^{0},-\vec{x})$, and vice versa. Consequently, the curvatures transform as $R^{(+)}_{ab}(x^{0},\vec{x})\mapsto R^{(-)}_{ab}(x^{0},-\vec{x})$ and $R^{(-)}_{ab}(x^{0},\vec{x})\mapsto R^{(+)}_{ab}(x^{0},-\vec{x})$. The scalar field $\phi^{ab}(x^{0},\vec{x})$ is assumed to be parity-even, transforming as $\phi^{ab}(x^{0},\vec{x})\mapsto\phi^{ab}(x^{0},-\vec{x})$.
Next note that the Einstein-Hilbert-like term is constructed as a sum over both chiral sectors and takes the form:
$\mathcal{L}_{\text{grav}}=\frac{1}{32\pi G}\,\epsilon^{abcd}\,e_{a}\wedge e_{b}\wedge\left(R^{(+)}_{cd}+R^{(-)}_{cd}\right),$
ensuring that parity symmetry is respected at the level of the classical action. The Yang-Mills-type kinetic terms are introduced independently for each chiral component and are written as:
$\mathcal{L}_{\text{YM}}=\frac{1}{4g_{+}^{2}}R^{(+)}_{ab}\wedge\star R^{ab(+)}+\frac{1}{4g_{-}^{2}}R^{(-)}_{ab}\wedge\star R^{ab(-)}.$
These terms preserve parity symmetry provided the chiral sectors are treated on equal footing.
To introduce the possibility of spontaneous parity violation, we couple an antisymmetric tensor scalar field $\phi^{ab}$ to the difference of the two chiral kinetic terms. This field transforms in the adjoint representation of the local Lorentz group and is assumed to be real. Its coupling to the curvature is written as
$\mathcal{L}_{\text{break}}=\frac{\alpha}{2}\,\phi^{bc}\left(R^{(+)}_{ab}\wedge\star R^{a(+)}_{~{}c}+R^{(-)}_{ab}\wedge\star R^{a(-)}_{~{}c}\right),$
where $\alpha$ is a dimensionless coupling constant and $\eta_{cd}$ is the Minkowski metric on the internal indices. This term is manifestly parity-invariant under the simultaneous exchange (+)()(+$\leftrightarrow$(-)( + ) ↔ ( - ) and the invariance of $\phi^{ab}$.
Finally, the dynamics of the scalar field is governed by a standard kinetic term and a Higgs-like potential. The kinetic term is constructed using the Lorentz-covariant derivative $D\phi^{ab}=d\phi^{ab}+\omega^{a}{}_{c}\phi^{cb}+\omega^{b}{}_{c}\phi^{ac}$, and takes the form $-\frac{1}{2}D\phi^{ab}\wedge*D\phi_{ab}$. The potential is given by
$V(\phi)=\frac{\lambda}{4}\left(\phi^{ab}\phi_{ab}-v^{2}\right)^{2},$
where $\lambda$ is a coupling constant and $v$ sets the symmetry-breaking scale. When $\phi^{ab}$ acquires a vacuum expectation value, this induces a spontaneous breaking of parity symmetry by generating unequal effective dynamics for the left- and right-handed chiral sectors.
@@
This theory is thus parity symmetric at the level of the action, and explicitly so before spontaneous symmetric breaking. However, once the scalar field condenses and selects a preferred direction in the internal space, the coupling between $\phi$ and the chiral curvatures breaks the parity symmetry spontaneously, leading to asymmetric effective dynamics in the gravitational sector. Minimizing the potential $V(\phi)$ gives a non-zero vacuum expectation value (vev) of the form:
$\langle\phi^{ab}\rangle=\frac{v}{2}\eta^{ab}$
Substituting this into the coupling term, together with the kinetic terms yields:
$\mathcal{L}_{\text{YM}}+\mathcal{L}_{\text{break}}=\frac{1}{4}\left(\frac{1}{g_{+}^{2}}+\alpha v\right)R^{(+)}_{ab}\wedge\star R^{ab(+)}+\frac{1}{4}\left(\frac{1}{g_{-}^{2}}+\alpha v\right)R^{(-)}_{ab}\wedge\star R^{ab(-)}.$ (6)
At low energies, if the symmetry-breaking scale satisfies
$v\gg\frac{1}{g_{-}^{2}}$
@@
the coefficient of the right handed kinetic term becomes very large, effectively suppressing it. As a result, the right-handed connection becomes non-dynamical and only the left-handed kinetic term remains:
$\mathcal{L}_{\text{YM}}^{\text{eff}}=\frac{1}{4g_{+}^{2}}R^{(+)}_{ab}\wedge\star R^{ab(+)},$ (7)
where we have redefined
$\frac{1}{g_{+}^{2}}+\alpha v\to\frac{1}{g_{+}^{2}}$
and so the left-handed coupling constant can be controlled by the Higgs vev.
@@
The fermion coupling is introduced through the chiral covariant derivative:
$\mathcal{L}_{\text{fermion}}=\frac{i}{2}\star e_{a}\wedge(\bar{\psi}_{L}\gamma^{a}D^{(+)}\psi_{L}+\bar{\psi}_{R}\gamma^{a}D^{(-)}\psi_{R}),$ (8)
@@
where the exterior covariant derivative acting on the spinors is defined by
$D^{(\pm)}\psi_{L/R}=d\psi_{L/R}+\frac{1}{4}\omega^{(\pm)ab}\gamma_{ab}\psi_{L/R}$ (9)
where $\psi_{L/R}=P_{L/R}\psi$ are the usual left- and right-handed spinor components and $\gamma_{ab}=\gamma_{[a}\gamma_{b]}$. After symmetry breaking, the right-handed connection becomes non-dynamical, reducing the fermion coupling to the left-handed chiral component:
$\mathcal{L}^{\text{eff}}_{\text{fermion}}=\frac{i}{2}\star e_{a}\wedge\bar{\psi}_{L}\gamma^{a}D^{(+)}\psi_{L}.$ (10)
@@
Hence, after symmetry breaking and adding the spinor term to (3) we obtain:
$\mathcal{L}^{\text{eff}}_{\text{total}}=\;$ $\frac{1}{32\pi G}\,\epsilon^{abcd}\,e_{a}\wedge e_{b}\wedge R^{(-)}_{cd}$
$+\frac{1}{4g_{+}^{2}}\bar{F}_{i}\wedge\star\bar{F}^{i}$
$+\frac{i}{2}\star e_{a}\wedge\bar{\psi}_{L}\gamma^{a}D^{(+)}\psi_{L},$ (11)
@@
which is of course standard gravity in the Ashtekar self-dual formulation plus weak theory, with fermions. We have relabeled the weak curvature as $\bar{F}^{i}=2iR^{(+)0i}$ to put it in a more familiar form. Notice we also see that the weak sector only couples to left handed spinors, reflecting chirality and parity violation.
@@
The equations of motion of our theory can be derived by the usual variation of the Lagrangian with respect to the fields, these being the tetrad $e^{a}$, both left- and right-handed connections $\omega^{ab(\pm)}$ and the spinor $\psi_{L}$. Hence, we get, respectively,
$\epsilon_{abcd}\left[\frac{1}{16\pi G}e^{b}\wedge R^{cd(-)}+\frac{i}{4}\bar{\psi}_{L}\gamma^{b}D^{(+)}\psi_{L}\wedge e^{c}\wedge e^{d}\right]=0$
$\star e_{a}\bar{\psi}_{L}\gamma^{a}\gamma_{cd}\psi_{L}=0$
$\epsilon_{abcd}T^{a(-)}\wedge e^{b}=0$
$\star e_{a}\gamma^{a}\wedge D^{(+)}\psi_{L}=0,$ (12)
@@
$\omega^{(+)}_{ab}=(\omega^{(+)}_{ab})^{*}$ (14)
where $\omega^{(-)}_{ab}(e)$ is the right-handed spin connection which possesses vanishing torsion: $\mathrm{d}\omega^{(-)}_{ab}(e)+\omega^{(-)}_{ab}(e)\wedge e^{b}=0$.

### v2

@@ I Introduction and Motivation
The unification of gravity with the other fundamental interactions remains one of the most profound challenges in theoretical physics. While the Standard Model successfully unifies the electroweak and strong interactions within a gauge-theoretic framework, gravity remains conceptually and technically distinct due to its geometric nature. An intriguing approach to unification is the idea that gravity and the weak interaction could emerge from a common underlying chiral gauge theory. This framework is referred to as graviweak unification Alexander:2012ge ; Alexander:0706.4481 ; Nesti:0706.3307 ; Nesti:0706.3304 . The motivation stems from the realization that both gravity and the weak interaction can be formulated as chiral theories. The weak interaction is intrinsically chiral, only left-handed fermions couple to the weak gauge bosons, while gravity, in the Einstein-Cartan formalism, can also be made chiral by splitting the spin connection into left- and right-handed components associated with the group structure KirillBook ; Freidel:2005mq ; Capovilla:1991qt ; Plebanski:1977pw ; Celada:2016jdt ; Alexandre:2407.19363
@@
These considerations motivate the construction of a theory in which gravity is initially described by both chiral components. A spontaneous symmetry-breaking mechanism then suppresses the right-handed chiral component at low energies. The left-handed component would naturally couple to left-handed fermions, mimicking the chiral nature of the weak interaction. In particular, the gravitational interaction should have the opposite chirality of the weak interaction, reflecting the parity-breaking structure of the Standard Model. This paper develops a model of graviweak unification where a Higgs-like field higgs generates this parity asymmetry dynamically through a $Z_{2}$ symmetric potential. The right-handed kinetic term is suppressed, leaving a chiral gravitational sector.
We develop these ideas in this paper along the following logical thread. In Section II we propose a gravitational theory formulated in terms of chiral components of the Lorentz connection and tetrad fields. The total Lagrangian includes Einstein-Hilbert-type terms, Yang-Mills-like kinetic terms for the chiral curvatures, and a Higgs-like pseudoscalar field $\phi$ that couples to the the parity odd terms and dynamically breaks parity symmetry through its vacuum expectation value. Thus, we have a theory that unifies gravity and the weak interaction in a parity invariant construction before symmetry breaking, reducing to the Ashtekar self-dual formulation of gravity plus the weak sector after symmetry breaking. This will be the foundations of this paper, around which we will investigate developments and variations.
In Section III we then rewrite this theory in the two-spinor notation, the natural language for treating gravity and chiral gauge fields on the same footing. This formalism is especially convenient for formulating unification frameworks such as graviweak models and it points us towards a new development: parity symmetry breaking that goes hand in hand with a larger gauge group. This receives further exploration in Section IV, where we extend the internal Lorentz group underpinning the basic construction to $SO(1,3+N)$. A major implication is the appearance of possible interactions between the gravitational 2-form curvature $R^{a}_{~b}$ of $SO(1,3)$ and the Yang-Mills field strength $F^{ij}$ of a $SO(N)$ group. This is a remarkable novelty with regards to previous graviweak models. We close this paper with a few speculations on the testability and falsifiability of this scenario.
@@ II The Graviweak-Higgs System
We start by laying down the structure of the theory, which is initially parity invariant and built from left $(+)$ and right-handed ()(-) chiral curvature 2-forms. This is done by introducing the projection operators:
[…]
@@
$P^{(\pm){cd}}{}_{ab}\;\omega_{cd}=\omega^{(\pm)}_{ab}=\frac{1}{2}\left(\omega_{ab}\pm\frac{i}{2}\epsilon_{ab}{}^{cd}\omega_{cd}\right),$ (2)
and hence the curvature 2-forms are:
$R^{(\pm)}_{ab}=d\omega^{(\pm)}_{ab}+\omega^{(\pm)}_{a}{}^{c}\wedge\omega^{(\pm)}_{cb}$. (3)
We note the following useful definitions of Hodge stars. The $\widetilde{\ }$ denotes the Hodge dual on the spacetime indices and the star symbol $\star$ will be used to denote the Hodge dual on the group indices:
$\widetilde{R}^{ab}_{\mathrm{\ \ }\mu\nu}=\frac{1}{2}\epsilon_{\rho\sigma\mu\nu}R^{ab\rho\sigma},$
$\star R^{ab}_{\mathrm{\ \ }\mu\nu}=\frac{1}{2}\epsilon^{abcd}R_{cd\mu\nu}.$
These definitions coupled with equation (2), show that applying the Hodge star in the group, to chiral terms yields:
${\star}R^{(-)}_{ab}\;=\;i\,R^{(-)}_{ab}$, (4)
${\star}R^{(+)}_{ab}\;=\;-\,i\,R^{(+)}_{ab}$. (5)
@@
With these in mind we then consider the most general parity-invariant gauge theory in $SO(1,3)$ containing only linear and quadratic terms in the curvature, with the condition that the theory should result in general relativity and the weak sector after spontaneous symmetry breaking. Its Lagrangian is
@@
$\mathcal{L}_{\text{total}}=\;$ $\frac{1}{16\pi G}\left[\,\,e^{a}\wedge e^{b}\wedge\star R_{ab}+i\phi\>\>e^{a}\wedge e^{b}\wedge R_{ab}\right]$
$-\frac{1}{8g^{2}}\left[R_{ab}\wedge\widetilde{R}^{ab}+i\phi\>R_{ab}\wedge\star\widetilde{R}^{ab}\right]$
$-\frac{1}{2}d\phi\wedge\widetilde{d}\phi-{V(\phi)}\widetilde{1}$. (6)
@@
A few comments on this action are in order. First note that the parity transformations under which this theory is initially symmetric are defined as follows. The tetrad transforms as $e^{a}(x^{0},\vec{x})\mapsto e^{a}(x^{0},-\vec{x})$, and the chiral components of the connection exchange under parity: $\omega^{(+)}_{ab}(x^{0},\vec{x})\mapsto\omega^{(-)}_{ab}(x^{0},-\vec{x})$, and vice versa. Consequently, the curvatures transform as $R^{(+)}_{ab}(x^{0},\vec{x})\mapsto R^{(-)}_{ab}(x^{0},-\vec{x})$ and $R^{(-)}_{ab}(x^{0},\vec{x})\mapsto R^{(+)}_{ab}(x^{0},-\vec{x})$. The pseudoscalar field $\phi(x^{0},\vec{x})$ is to be parity-odd, transforming as $\phi(x^{0},\vec{x})\mapsto-\phi(x^{0},-\vec{x})$.
We consider a modification of the chiral gravitational action in which the couplings to parity odd terms such as the Holst term are modified to parity even by the pseudoscalar field $\phi$. This ensures that parity symmetry is respected at the level of the classical action. The Lagrangian is constructed to be explicitly parity-invariant, even while allowing for the spontaneous breaking of parity through the vacuum structure of $\phi$: $V(\phi)$ is a parity-invariant potential, of the Higgs form
$V(\phi)=\frac{\lambda}{4}(\phi^{2}-1)^{2},$
leading to spontaneous symmetry breaking with vacuum expectation values (vev) $\langle\phi\rangle=\pm 1$ and where $\lambda$ is a coupling constant. Under a parity transformation, the exchange of the parity odd terms is exactly compensated by $\phi\mapsto-\phi$, ensuring that the Lagrangian remains invariant. Once $\phi$ acquires a nonzero vacuum expectation value, this symmetry is spontaneously broken. The couplings induce a differential scaling between the self-dual and anti-self-dual sectors. For example, for $\langle\phi\rangle>0$, one sees, once factoring out $i$ that the first line of the Lagrangian becomes the right handed projection operator in (1) times a Holst term: $\frac{i}{8\pi G}\;e^{a}\wedge e^{b}\wedge P^{(-)cd}{}_{ab}R^{cd}=\frac{i}{8\pi G}\;e^{a}\wedge e^{b}\wedge R^{(-)}_{ab}$; for $\langle\phi\rangle<0$, the opposite occurs $\frac{i}{8\pi G}\;e^{a}\wedge e^{b}\wedge P^{(+)cd}{}_{ab}R^{cd}=\frac{i}{8\pi G}\;e^{a}\wedge e^{b}\wedge R^{(+)}_{ab}$. This leads to dynamical chirality selection, where the vacuum structure of the pseudoscalar field determines which chiral sector governs the low-energy gravitational dynamics.
The field $\phi$ in this framework behaves as a chiral modulus: it continuously controls the binary weighting of curvature contributions from the two chiral sectors. This construction introduces a novel, geometrically natural framework in which parity remains a fundamental symmetry of the action, while chirality can be selected dynamically through spontaneous symmetry breaking. It defines a new class of moduli-driven parity-variant theories of gravity within a chiral, pseudoscalar, and parity-respecting structure.
@@
This theory is thus parity symmetric at the level of the action, and explicitly so before spontaneous symmetry breaking. However, once the pseudoscalar field condensates, the coupling between $\phi$ and the chiral curvatures breaks the parity symmetry spontaneously, leading to asymmetric effective dynamics in the gravitational sector. Substituting the vev into (6) and employing equations (1) and (4) yields:
$\mathcal{L}=-\frac{1}{4g^{2}}R^{(+)}_{ab}\wedge\widetilde{R}^{ab(+)}+\frac{1}{8\pi G}\;e^{a}\wedge e^{b}\wedge\star R^{(-)}_{ab}$.
@@
As a result, the right-handed connection becomes purely gravitational, meaning it is only present in the Lagrangian by coupling to the tetrad, and only the left-handed kinetic term remains:
$\mathcal{L}_{\text{YM}}^{\text{eff}}=-\frac{1}{4g^{2}}R^{(+)}_{ab}\wedge\widetilde{R}^{ab(+)}.$ (7)
@@
The parity even coupling of the fermions is introduced through the general covariant derivative:
$\mathcal{L}_{\text{fermion}}=-\frac{i}{2}\star e_{a}\wedge\bar{\psi}\gamma^{a}D\psi.$ (8)
@@
The spin connection can then be decomposed into chiral parts, and thus the exterior covariant derivative acting on the spinors is defined by
$D^{(\pm)}\psi_{L/R}=d\psi_{L/R}+\frac{1}{4}\omega^{(\pm)ab}\gamma_{ab}\psi_{L/R},$ (9)
where $\psi_{L/R}=P_{L/R}\psi$ are the usual left- and right-handed spinor components and $\gamma_{ab}=\gamma_{[a}\gamma_{b]}$. It can be shown that the product of projection operators of the opposing chirality between $P_{L/R}$ and $P^{(\pm)}$ is zero and such the fermion Lagrangian reduces to:
$\mathcal{L}_{\text{fermion}}=-\frac{i}{2}\star e_{a}\wedge[\bar{\psi}_{L}\gamma^{a}D^{(+)}\psi_{L}+\bar{\psi}_{R}\gamma^{a}D^{(-)}\psi_{R}].$ (10)
@@
Hence, after symmetry breaking and adding the spinor terms to (6) we obtain:
$\mathcal{L}^{\text{eff}}_{\text{total}}=\;$ $\frac{1}{16\pi G}\,\epsilon^{abcd}\,e_{a}\wedge e_{b}\wedge R^{(-)}_{cd}$
$-\frac{1}{4g^{2}}\bar{F}_{i}\wedge\widetilde{\bar{F}}^{i}$
$-\frac{i}{2}\star e_{a}\wedge[\bar{\psi}_{L}\gamma^{a}D^{(+)}\psi_{L}+\bar{\psi}_{R}\gamma^{a}D^{(-)}\psi_{R}],$ (11)
@@
which is of course standard gravity in the Ashtekar self-dual formulation plus weak theory, with fermions. We have relabeled the weak curvature as $\bar{F}^{i}=2iR^{(+)0i}$ to put it in a more familiar form. Notice we also see that the weak sector only couples to left handed spinors, reflecting chirality and parity violation, and the right-handed ones only couple to gravity.
@@
The equations of motion of our theory can be derived by the usual variation of the Lagrangian with respect to the fields, these being the tetrad $e^{a}$, both left- and right-handed connections $\omega^{ab(\pm)}$ and the spinors $\psi_{L/R}$. Hence, we get, respectively,
$\epsilon_{abcd}\Bigg[\frac{1}{8\pi G}e^{b}\wedge R^{cd(-)}+$
$+\frac{i}{4}\bar{\psi}_{L}\gamma^{b}D^{(+)}\psi_{L}\wedge e^{c}\wedge e^{d}+\frac{i}{4}\bar{\psi}_{R}\gamma^{b}D^{(-)}\psi_{R}\wedge e^{c}\wedge e^{d}\Bigg]=$
$=\frac{1}{4g^{2}}\left[R^{(+)cd}(\widetilde{R}^{(+)}_{cd})_{ab}-(R^{(+)cd})_{ab}\widetilde{R}^{(+)}_{cd}\right]e^{b}$
$D^{(+)}\widetilde{R}^{(+)}_{ab}=\frac{ig^{2}}{4}\star e_{c}\bar{\psi}_{L}\gamma^{c}\gamma_{ab}\psi_{L}$
$\epsilon_{abcd}T^{(-)c}\wedge e^{d}=2\pi Gi\star e_{c}\wedge\bar{\psi}_{R}\gamma^{c}\gamma_{ab}\psi_{R}$
$\star e_{a}\gamma^{a}\wedge D^{(+)}\psi_{L}=0,$
$\star e_{a}\gamma^{a}\wedge D^{(-)}\psi_{R}=0,$ (12)
@@
$\bar{A}_{i}=(\bar{A}_{i})^{*}$ (14)
where $\omega^{(-)}_{ab}(e)$ is the right-handed spin connection which possesses vanishing torsion: $\mathrm{d}\omega^{(-)}_{ab}(e)+\omega^{(-)}_{ab}(e)\wedge e^{b}=0$. Furthermore, $\bar{A}_{i}=2i\omega^{(+)}_{0i}$ is the $SU(2)$ gauge field for $\bar{F}_{i}$:
$\bar{F}_{i}=d\bar{A}_{i}+\frac{1}{2}\epsilon_{ijk}\bar{A}^{j}\wedge\bar{A}^{k}.$ (15)